OpenStudy (anonymous):

i need some help from any one of you dudes on this site! girl in distress!math sucks:(

OpenStudy (anonymous):

Do you need a math tutor or something ?

OpenStudy (anonymous):

no i just need help with these questions then we wont ever have to talk again promise:)

OpenStudy (anonymous):

mystery questions?

OpenStudy (anonymous):

the first is The following are the lengths of two sides of a right triangle: one leg is 3 cm and the hypotenuse is 5 cm. Find the length of the missing leg.

OpenStudy (anonymous):

idk!!!!

OpenStudy (anonymous):

mother of all right triangles

OpenStudy (anonymous):

lol:)

OpenStudy (anonymous):

most famous of all famous right triangles i bet you know it

OpenStudy (anonymous):



OpenStudy (anonymous):

\[3^2+b^2=5^2\] solve for \(b\)

OpenStudy (the_fizicx99):

Hypotenuse formula is: a^2 + b^2 = c^2 Where A or B are on of the legs, C is the hypotenuse. Let's call a 3, 3^2 + b^2 = 5^2 9 + b^2 = 25 -9 -9 b^2 = 16 √b^2 = √16 b = 4 3^2 + 4^2 = 5^2 9 + 16 = 25 25 = 25

OpenStudy (the_fizicx99):

one* U_U

OpenStudy (anonymous):

wow you guys rock! all of you:) so #2..

OpenStudy (anonymous):

wait it cant be 25

OpenStudy (anonymous):



OpenStudy (the_fizicx99):

Wot? 5 is the hypotenuse, 5^2 = 25 xD

OpenStudy (acxbox22):

yeah, the most basic pythagorean triple

OpenStudy (anonymous):

but it can be four! ya! thank you!

OpenStudy (the_fizicx99):

.-. B is 4 X_X A = 3 B = 4 C = 5 >.>

OpenStudy (anonymous):

#2 is The following are the lengths of the legs of a right triangle: 12 yd and 9 yd. Find the missing side.

OpenStudy (anonymous):

same triangle multiply each by 3

OpenStudy (anonymous):

\[3,4,5\\ 3\times 3,4\times 3,5\times 3\\ 9,12,15\]

OpenStudy (the_fizicx99):

True ^, but it's easier to see which leg is which >.<

OpenStudy (anonymous):

it is?

OpenStudy (the_fizicx99):

Don't question satellite >.>

OpenStudy (anonymous):

yeah, i have no answers anyways

OpenStudy (anonymous):

im not questioning anyone im just really confused!

OpenStudy (anonymous):

and stupide

OpenStudy (the_fizicx99):

xD well see that the new sides are 12 and 9? 4 * 3 = 12, 3 * 3 = 9 So with logic you can assume c = 15, since c = 5, 5 * 3 = 15

OpenStudy (anonymous):

15... ya ok that works. thanks for showing your work! and by the way i have 10 questions left

OpenStudy (the_fizicx99):

o_o Have fun! >.>

OpenStudy (anonymous):

thanks! lol:)

OpenStudy (the_fizicx99):

;D

OpenStudy (anonymous):

so are you really going to leave me hangin?

OpenStudy (anonymous):

i just need someone to get me through these! its late

OpenStudy (anonymous):

wtb2tbmkb1opmgrpomrbeqrberbqerb3rb

OpenStudy (anonymous):

need more tutoring ?

OpenStudy (anonymous):

yes please!

OpenStudy (anonymous):

So, you know how to find the hypotenuse or any other side of the right triangle, right ?

OpenStudy (anonymous):

yes but the questions are different now!!!they changed:(

OpenStudy (anonymous):

which questions ? which problems are you doing right now ?

OpenStudy (anonymous):

prob. 16. -7 +(_4)

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

is it, number 16. -7+(-4) ?

OpenStudy (anonymous):

yes! :( im so dumb:( please help

OpenStudy (anonymous):

when you say " b+ (-a) " it is same as just saying " b - a " . So -7+(-4) = -7 - 4 = you tell me. Can you ?

OpenStudy (anonymous):

ya its -11 right?

OpenStudy (anonymous):

YES !

OpenStudy (anonymous):

sweet!!!!!! one down for me! you rock:)

OpenStudy (anonymous):

tnx :)

OpenStudy (anonymous):

no prob.Hey your picture changed!

OpenStudy (anonymous):

sorry i should focuse on math.

OpenStudy (anonymous):

it's ok... I changed it just now.